\begin{thebibliography}{99}

\bibitem{chen-zhang}Z. Chen, W. Lin, Square-mean weighted pseudo almost automorphic solutions for non-autonomous stochastic evolution equations, J. Math. Pures Appl. 100 (2013), 476--504.


\bibitem{blot-waa1} J. Blot, G. M. Mophoub, G. M. N'Gu\'{e}r\'{e}kata, D. Pennequina, Weighted pseudo almost automorphic functions and applications n to abstract differential equations, Nonlinear Anal. TMA 71 (2009), 903--909.

\bibitem{blot-waa2}J. Blot, P. Cieutat, K. Ezzinbi, Measure theory and pseudo almost automorphic functions:
new developments and applications, Nonlinear Anal. TMA 75 (2012),
2426--2447.


\bibitem{bochner}S. Bochner, A new approach to almost-periodicity, Proc. Natl. Acad. Sci. USA 48 (1962), 2039--2043.


\bibitem{diagana-wpap}T. Diagana, Weighted pseudo almost periodic functions and
applications, C. R. Acad. Sci. Paris, Ser. I 343 (10) (2006),
643--646.

\bibitem{ding}H. S. Ding, W. long, G. M. N'Gu\'{e}r\'{e}kata, A composition theorem for weighted pseudo-almost automorphic functions and applications, Nonlinear Anal. TMA 73 (2010), 2644--2650.


\bibitem{ding2014}H. S. Ding, J. Liang, T. J. Xiao, Weighted pseudo almost automorphic functions and WPAA solutions
to semilinear evolution equations, J. Math. Anal. Appl. 409 (2014),
409--427.


\bibitem{zhang}D. Ji, C. Zhang, Translation invariance of weighted
pseudo almost periodic functions and related problems, J. Math.
Anal. Appl. 391 (2012), 350--362.


\bibitem{liu}J. Liu, X. Song, Almost automorphic and weighted
pseudo almost automorphic solutions of semilinear evolution
equations, J. Funct. Anal. 258 (2010), 196--207.






\bibitem{gaston05}G. M. N'Gu\'{e}r\'{e}kata, Topics in almost
automorphy, Springer-Verlag, New York, 2005.





\bibitem{zhang-book}C. Zhang, Almost Periodic Type Functions and Ergodicity, Kluwer
Academic Publishers, Dordrecht, Boston, London, 2003.






%\bibitem{n1}S. Abbas, Pseudo almost automorphic solutions of some nonlinear integro-differential equations, Comput. Math. Appl. 62 (2011), 2259--2272.

%\bibitem{diagana-book}T. Diagana, Pseudo almost periodic functions in Banach spaces, Nova Science Publishers, Inc., New York, 2007.

%\bibitem{diagana2011}T. Diagana, The existence of a weighted mean for almost periodic functions, Nonlinear Anal. 74 (2011), 4267--4273.


%\bibitem{blot-waa1}J. Blot, P. Cieutat, K. Ezzinbi, New approach for weighted pseudo almost periodic functions
%under the light of measure theory, basic results and applications,
%Appl. Anal. 92 (2013), 493--526.



%\bibitem{corduneanu}C. Corduneanu, Almost Periodic Functions, 2nd edition, Chelsea, New york, 1989.



%\bibitem{ding1}H. S. Ding, G. M. N'Gu\'{e}r\'{e}kata, J. J. Nieto, Weighted pseudo almost periodic solutions for a class of discrete hematopoiesis model, Revista Matematica Complutense, 26 (2013), 427--443.






%\bibitem{ding2013}H. S. Ding, J. Liang, T. J. Xiao, Weighted pseudo almost periodic functions and applications to evolution equations with delay, Appl. Math. Comput. 219 (2013), 8949--8958.





%\bibitem{liang}J. Liang, T. J. Xiao, J. Zhang, Decomposition of weighted pseudo-almost periodic functions, Nonlinear Anal. TMA 73 (2010), 3456--3461.





%\bibitem{zhang1994}C. Zhang, Pseudo almost periodic solutions of some differential equations, J. Math. Anal. Appl. 181 (1994), 62--76.


%\bibitem{zhang1997}C. Zhang, Vector-valued pseudo almost periodic functions, Czech. Math. J. 47(122) (1997), 385--394.




%\bibitem{jun-zhang}J. Zhang, T. J. Xiao, J. Liang, Weighted pseudo almost periodic functions and applications to semilinear evolution equations, Abst. Appl. Anal. 2012 (2012) ArticleID 179525.






\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{amsmath}
\usepackage{cases}
\usepackage{amssymb}
\usepackage{amsfonts}

\usepackage{bbm}
\usepackage{amsmath,amsthm,amssymb,amscd}
\usepackage{latexsym}
\usepackage{hyperref}
\usepackage[numbers,sort&compress]{natbib}
\usepackage{hypernat}
\allowdisplaybreaks
\def\r{\mathbb{R}}

\def\n{\mathbb{N}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{definition}[theorem]{Definition}

\theoremstyle{definition} \theoremstyle{remark}
\numberwithin{equation}{section}
\begin{document}
%------------------------------------------------------------------------

\title{{\bf On completeness of the space of weighted pseudo almost automorphic functions\thanks{The work was supported by
the NSF of China (11461034), and the Program for Cultivating Young
Scientist of Jiangxi Province (20133BCB23009).}}}
\date{}
\author{Zhe-Ming Zheng, Hui-Sheng Ding\thanks{Corresponding author. E-mail address:
dinghs@mail.ustc.edu.cn (H.-S. Ding)} \\
{\small { \it College of Mathematics and Information Science,
Jiangxi Normal University}}\\{\small {\it Nanchang, Jiangxi 330022,
People's Republic of China }}}
\maketitle

\begin{abstract}
In this paper, we prove that for every $\rho\in \mathbb{U}_{\infty}$, the space of weighted pseudo almost automorphic functions is complete under the supremum norm. This gives an affirmative answer to a key and fundamental problem for weighted pseudo almost automorphic functions, and fills a gap in the proof of [J. Funct. Anal. 258, No. 1, 196-207 (2010)].

\textbf{Keywords:} complete, weighted pseudo almost automorphic, almost automorphic.

\textbf{2000 Mathematics Subject Classification:} 43A60.
\end{abstract}

\section{Introduction}

The notion of weighted pseudo almost automorphic functions is initiated
by Blot et al. \cite{blot-waa1}, which is an interesting
generalization of the classical almost automorphic functions introduced by Bochner \cite{bochner}, as
well as a generalization of the notion of weighted pseudo almost periodic
function introduced by Diagana \cite{diagana-wpap}.

Since the work of Blot et al. \cite{blot-waa1}, there has been of great interest for many
authors to investigate weighted pseudo almost automorphic functions and their applications to evolution equations (see, e.g., \cite{liu,chen-zhang,blot-waa2,ding2014,zhang}).

Especially, in \cite{liu}, the authors discuss the existence and uniqueness of weighted pseudo almost automorphic mild solution to
a class of semilinear evolution equation
$$x'(t)=A(t)x(t)+f(t,x(t))$$
in a Banach space. However, just as noted in Zbl 1194.47047, in the proof of \cite[Theorem 4.2]{liu}, the authors use the Banach contraction mapping principle, and thus, the completeness of the space of weighted pseudo almost automorphic functions is needed. In fact, to the best of our knowledge, all the results concerning the completeness of the space of weighted pseudo almost automorphic functions need some restrictive conditions on the weighted term $\rho$ (cf. \cite{blot-waa2,ding2014,zhang}). \textit{There is no proof} for the completeness of the space of weighted pseudo almost automorphic functions with no restrictive conditions on $\rho$ until now.

Thus, just as indicated in Zbl 1194.47047, there is a gap in the proof of \cite[Theorem 4.2]{liu}. In addition, the completeness of the space of weighted pseudo almost automorphic functions is very important for applications of such functions. However, due to the influence of
weighted term, this problem becomes very tricky. We refer the reader to \cite{blot-waa2,ding2014,zhang} for some recent results on the completeness of the space of weighted pseudo almost automorphic functions. But, as we noted in the above paragraph, all the earlier results need some restrictive conditions on the weighted term $\rho$.

In this paper, we aim to fill this gap in the proof of \cite[Theorem 4.2]{liu} and solve this key problem completely.

Throughout the rest of this paper, we denote by $\n$ the set of
positive integers, by $\mathbb{R}$ the set of real numbers, by $X$ a
Banach space, by $BC(\mathbb{R},X)$ the Banach space of all bounded
continuous functions $f:\r\to X$ with supremum norm, and by $\mathbb{U}$ the set of
functions (weights) $\rho:\mathbb{R}\to[0,+\infty),$ which is
locally integrable over $\mathbb{R}$. In addition, for $\rho\in
\mathbb{U}$ and $r>0$, we denote
$$
\mu(r,\rho):=\int_{-r}^{r}\rho(t)dt.$$
$$
\mathbb{U}_{\infty}:=\{\rho\in\mathbb{U}:\lim_{r\to+\infty}\mu(r,\rho)=+\infty\},
$$
and
$$
\mathbb{U}_{B}:=\{\rho\in\mathbb{U}_{\infty}:\rho \text{ is bounded
with }  \inf_{t\in\mathbb{R}}\rho(t)>0\}.
$$
Obviously,
$\mathbb{U}_{B}\subset\mathbb{U}_{\infty}\subset\mathbb{U}$, with
strict inclusions.

Next, let us recall some notions and basic results about almost automorphic type functions and almost periodic type functions (for more
details, see \cite{blot-waa1,zhang-book,diagana-wpap,gaston05}).

\begin{definition}A set $P \subset\r$ is called relatively dense in $\r$ if there exists a number $l>0$
such that $\forall a\in \mathbb{R}$, $[a,a+l]\bigcap
P\neq\varnothing.$
\end{definition}

\begin{definition} A continuous function $f:\r\to X$ is called almost periodic if for
every $\varepsilon>0$ there exists a relatively dense set
$P(\varepsilon,f)$ such that
$$\sup_{t\in\r}\|f(t+\tau)-f(t)\|<\varepsilon$$
for all $\tau\in P(\varepsilon,f)$. We denote the set of all such
functions by $AP(X)$.
\end{definition}

\begin{definition}\label{aa}
A continuous function $f:\mathbb{R}\rightarrow X$ is called almost
automorphic if for every real sequence $(s_{m})$, there exists a
subsequence $(s_{n})$ such that
$$g(t):=\lim_{n\rightarrow\infty}f(t+s_{n})$$
is well defined for each $t\in \mathbb{R}$ and
$$\lim_{n\rightarrow\infty}g(t-s_{n})=f(t)$$
for each $t\in \mathbb{R}$. Denote by $AA(X)$ the set of all such
functions.
\end{definition}

For each $\rho\in\mathbb{U}_{\infty}$, define
$$
M_{0}(X,\rho):=\{f\in
BC(\mathbb{R},X):\lim_{r\to+\infty}\frac{1}{\mu(r,\rho)}
\int_{-r}^{r}\|f(t)\|\rho(t)dt=0\}.
$$

\begin{definition}
 Let $\rho\in\mathbb{U}_{\infty}$. A
function $f\in BC(\mathbb{R},X)$ \ is called weighted pseudo almost
automorphic or $\rho$-pseudo almost automorphic if it can be
expressed as $f=g+h$, where $g\in AA(X)$ and $h \in M_{0}(X,\rho)$.
The set of such functions will be denoted by $PAA(X,\rho)$.
\end{definition}

\begin{definition}
 Let $\rho\in\mathbb{U}_{\infty}$. A
function $f\in BC(\mathbb{R},X)$ \ is called weighted pseudo almost
periodic or $\rho$-pseudo almost periodic if it can be expressed as
$f=g+h$, where $g\in AP(X)$ and $h \in M_{0}(X,\rho)$. The set of
such functions will be denoted by $PAP(X,\rho)$.
\end{definition}

\begin{definition}{\rm \cite{zhang}}
Let $\rho\in\mathbb{U}_{\infty}$. A set $C\subset\r$ is said to be a
$\rho$-ergodic zero set if
$$\lim_{r\rightarrow{+\infty}}\frac{\int_{[-r,r]\cap C}\rho(t)dt}{\mu(r,\rho)}=0.$$
\end{definition}

The following Lemma is due to \cite[Lemma 3.2]{ding} (see also
\cite{zhang}).

\begin{lemma}\label{lemma1}Let $f\in BC(\r,X)$ and $\rho\in \mathbb{U}_{\infty}$. Then $f\in M_{0}(X,\rho)$ if and only if for
every $\varepsilon>0$, $M_{\varepsilon}(f)$ is a $\rho$-ergodic zero
set, where $M_{\varepsilon}(f):=\{t\in\r:\|f(t)\|\geq
\varepsilon\}$.
\end{lemma}

\section{Completeness of $PAA(X,\rho)$}

\begin{theorem}\label{key-theorem} For every $f\in PAA(X,\rho)$, there exists a decomposition $f=g_0+h_0$, where $g_0\in AA(X)$ and $h_0\in M_0(X,\rho)$, such that
$$\|g_0\|\leq \|f\|.$$
\end{theorem}
\begin{proof}Let $f=g+h$, where $g\in AA(X)$ and $h\in M_0(X,\rho)$. Set
$$C_n=\{t\in\r: \|g(t)\|> \|f\|+\frac{1}{n}\},\quad n\in\n.$$
Noting that $C_n\subset \{t\in\r: \|h(t)\|\geq \frac{1}{n}\}$, it follows from Lemma \ref{lemma1} that every $C_n$ is a $\rho$-ergodic zero
set. Let
$$g_n(t)=\left\{\begin{array}{c@{\quad,\quad }l}
g(t) & t\notin C_n,\\[0.36cm]
(\|f\|+\frac{1}{n})\cdot \frac{g(t)}{\|g(t)\|} & t\in C_n.
\end{array}\right. $$
Next, we show that $g_n$ have the following three properties:

I. For every $n\in\n$, $g_n$ is continuous on $\r$.

Let $t_0\in \r$ and $t_k\to t_0$. We divide the proof of $g_n(t_k)\to g_n(t_0)$ into three cases.

Case (i): if $\|g(t_0)\|>\|f\|+\frac{1}{n}$, then $\|g(t_k)\|>\|f\|+\frac{1}{n}$ for sufficiently large $k$ since $g$ is continuous, and thus $t_k\in C_n$ for sufficiently large $k$. Also, in this case, $t_0\in C_n$. Then, it follows that
$$\|g_n(t_k)- g_n(t_0)\|=(\|f\|+\frac{1}{n})\cdot \left\|\frac{g(t_k)}{\|g(t_k)\|} - \frac{g(t_0)}{\|g(t_0)\|} \right\|\to 0.$$

Case (ii): if $\|g(t_0)\|<\|f\|+\frac{1}{n}$, then $\|g(t_k)\|<\|f\|+\frac{1}{n}$ for sufficiently large $k$ since $g$ is continuous, and thus $t_k\notin C_n$ for sufficiently large $k$. Also, in this case, $t_0\notin C_n$. Then, it follows that
$$\|g_n(t_k)- g_n(t_0)\|=\|g(t_k)- g(t_0)\|\to 0.$$

Case (iii): if $\|g(t_0)\|=\|f\|+\frac{1}{n}$, then $t_0\notin C_n$. Noting that $\|g(t_k)\|\to\|f\|+\frac{1}{n}$, we have
$$\|g_n(t_k)- g_n(t_0)\|=\|g(t_k)- g(t_0)\|\cdot[1-\chi _{C_n}(t_k)]+\left\|(\|f\|+\frac{1}{n})\cdot \frac{g(t_k)}{\|g(t_k)\|}-g(t_0)\right\|\cdot \chi_{C_n}(t_k)\to 0,$$
where $\chi_{C_n}$ is the characteristic function on $C_n$, i.e., $\chi_{C_n}(t)=1$ for all $t\in C_n$ and $\chi_{C_n}(t)=0$ for all $t\notin C_n$.

II. For every $n\in\n$, $g_n\in AA(X)$.

Taking an arbitrary sequence $s'_k$, there exist a subsequence $s_k$ and a function $\overline{g}:\r\to X$ such that $g(t+s_k)\to \overline{g}(t)$ and
$\overline{g}(t-s_k)\to g(t)$ for all $t\in\r$. Next, for every fixed $t\in\r$, we show that $g_n(t+s_k)\to \overline{g}_n(t)$, where
$$\overline{g}_n(t)=\left\{\begin{array}{c@{\quad,\quad }l}
\overline{g}(t) & \|\overline{g}(t)\|\leq \|f\|+\frac{1}{n},\\[0.36cm]
(\|f\|+\frac{1}{n})\cdot \frac{\overline{g}(t)}{\|\overline{g}(t)\|} & \|\overline{g}(t)\|>\|f\|+\frac{1}{n}.
\end{array}\right. $$
If $\|\overline{g}(t)\|>\|f\|+\frac{1}{n}$, then for sufficiently large $k$, $\|g(t+s_k)\|>\|f\|+\frac{1}{n}$ and thus $t+s_k\in C_n$. Then, it follows that
$$g_n(t+s_k)=(\|f\|+\frac{1}{n})\cdot \frac{g(t+s_k)}{\|g(t+s_k)\|}\to (\|f\|+\frac{1}{n})\cdot \frac{\overline{g}(t)}{\|\overline{g}(t)\|}=\overline{g}_n(t).$$
If $\|\overline{g}(t)\|<\|f\|+\frac{1}{n}$, then for sufficiently large $k$, $\|g(t+s_k)\|<\|f\|+\frac{1}{n}$ and thus $t+s_k\notin C_n$. Then, it is easy to see that
$$g_n(t+s_k)=g(t+s_k)\to \overline{g}(t)=\overline{g}_n(t).$$
If $\|\overline{g}(t)\|=\|f\|+\frac{1}{n}$, then $\overline{g}_n(t)=\overline{g}(t)$, and
\begin{eqnarray*}
&&\|g_n(t+s_k)-\overline{g}_n(t)\|\\&=& \|g_n(t+s_k)-\overline{g}(t)\|\\
&=& \|g(t+s_k)-\overline{g}(t)\|\cdot[1-\chi_{C_n}(t+s_k)]+\left\|(\|f\|+\frac{1}{n})\cdot \frac{g(t+s_k)}{\|g(t+s_k)\|} -\overline{g}(t)\right\|\cdot \chi_{C_n}(t+s_k)\\
&\to& 0,
\end{eqnarray*}
i.e., $g_n(t+s_k)\to\overline{g}_n(t)$. By a similar proof, one can also show that $\overline{g}_n(t-s_k)\to g_n(t)$.

III. $g_n(t)$ is uniformly convergent on $\r$.

For all $n>m$, noting $C_m\subset C_n$, we have
$$\|g_n(t)-g_m(t)\|=\left\{\begin{array}{c@{\quad,\quad }l}
0 & t\notin C_n,\\[0.36cm]
\frac{1}{m}-\frac{1}{n}<\frac{1}{m} & t\in C_m,\\[0.36cm]
\left|\|f\|+\frac{1}{n}-\|g(t)\|\right|\leq \frac{1}{m}-\frac{1}{n}<\frac{1}{m} & t\in C_n,\ t\notin C_m.
\end{array}\right. $$
So $\{g_n\}$ is a Cauchy sequence in $BC(\r,X)$.

Let $g_0=\lim_{n\to\infty}g_n$. Then it follows from I-III that $g_0\in AA(X)$ since $AA(X)$ is a closed subspace of $BC(\r,X)$. Moreover, noting that $\|g_n\|\leq \|f\|+\frac{1}{n}$, we have
$\|g_0\|\leq \|f\|$. Let $h_n=f-g_n$ and $h_0=\lim_{n\to\infty}h_n$. Then $f=g_0+h_0$. Moreover, for every fixed $n\in\n$, $h_n\in BC(\r,X)$ and $h_n(t)=h(t)$ for all $t\notin C_n$. Then, for every $\varepsilon>0$
$$\{t\in\r:\|h_n(t)\|\geq \varepsilon\}\subset \{t\in\r:\|h(t)\|\geq \varepsilon\}\cup C_n $$
Noting that $\{t\in\r:\|h(t)\|\geq \varepsilon\}$ and $C_n$ are both $\rho$-ergodic zero sets, we conclude that $\{t\in\r:\|h_n(t)\|\geq \varepsilon\}$ is also a $\rho$-ergodic zero set, and thus by Lemma \ref{lemma1}, $h_n\in M_0(X,\rho)$. Then, we have $h_0\in M_0(X,\rho)$ since $M_0(X)$ is also a closed subspace of $BC(\r,X)$. This completes the proof.
\end{proof}

\begin{corollary}\label{ap}For every $f\in PAP(X,\rho)$, there exists a decomposition $f=g_0+h_0$, where $g_0\in AP(X)$ and $h_0\in M_0(X,\rho)$, such that
$\|g_0\|\leq \|f\|.$
\end{corollary}
\begin{proof}We only need to modify the proof of II in Theorem \ref{key-theorem}. In fact, it suffices to show that for every $\varepsilon>0$,
$$\|g_n(t+\tau)-g_n(t)\|<\varepsilon,\quad t\in\r,\ \tau\in P(\varepsilon,g).$$
Fix $n\in\n$. For all $t,t+\tau\notin C_n$, it is easy to see that $$\|g_n(t+\tau)-g_n(t)\|=\|g(t+\tau)-g(t)\|<\varepsilon.$$
For all $t,t+\tau\in C_n$, we have
$$\|g_n(t+\tau)-g_n(t)\|\leq \frac{\Big\|g(t+\tau)\cdot \|g(t)\|-\|g(t+\tau)\|\cdot g(t)\Big\|}{\|g(t)\|}\leq 2\|g(t+\tau)-g(t)\|<2\varepsilon.$$
For all $t\notin C_n$ and $t+\tau\in C_n$, we have
$$\|g(t)\|\leq \|f\|+\frac{1}{n}<\|g(t+\tau)\|,$$
and
\begin{eqnarray*}
\|g_n(t+\tau)-g_n(t)\|&=& \Big\| \frac{\|f\|+\frac{1}{n}}{\|g(t+\tau)\|}\cdot g(t+\tau)-g(t)\Big\|\\
&\leq& \|g(t+\tau)-g(t)\|+\Big|\|f\|+\frac{1}{n}-\|g(t+\tau)\|\Big|\\
&\leq& 2\|g(t+\tau)-g(t)\|<2\varepsilon.
\end{eqnarray*}
For all $t\in C_n$ and $t+\tau\notin C_n$, the proof is similar.
\end{proof}

\begin{theorem}For every $\rho\in \mathbb{U}_{\infty}$, $PAA(X,\rho)$ is a Banach space under the supremum norm.
\end{theorem}
\begin{proof}Let $\{f_n\}$ be a Cauchy sequence in $PAA(X,\rho)$. Then, we can choose a subsequence $\{f_{n_k}\}$ such that
$$\|f_{n_{k+1}}-f_{n_k}\|\leq \frac{1}{2^k}.$$
Let $f_{n_1}=g_{n_1}+h_{n_1}$, where $g_{n_1}\in AA(X)$ and $h_{n_1}\in M_0(X,\rho)$. By applying Theorem \ref{key-theorem} on $f_{n_2}-f_{n_1}$, there exist $g_0\in AA(X) $ and $h_0\in  M_0(X,\rho)$ such that $f_{n_2}-f_{n_1}=g_0+h_0$ and
$$\|g_0\|\leq \|f_{n_2}-f_{n_1}\|.$$
Let $g_{n_2}=g_0+g_{n_1}$ and $h_{n_2}=h_0+h_{n_1}$. Then $g_{n_2}\in AA(X)$, $h_{n_2}\in M_0(X,\rho)$, $f_{n_2}=g_{n_2}+h_{n_2}$ and
$$\|g_{n_2}-g_{n_1}\|=\|g_0\|\leq \|f_{n_2}-f_{n_1}\|\leq \frac{1}{2}.$$
Continuing by this way, one can get two sequences $\{g_{n_k}\}\subset AA(X)$ and $\{h_{n_k}\}\subset M_0(X,\rho)$ such that for all $k\in\n$, $f_{n_k}=g_{n_k}+h_{n_k}$ and
$$\|g_{n_{k+1}}-g_{n_k}\|\leq \|f_{n_{k+1}}-f_{n_k}\|\leq \frac{1}{2^k}.$$
Thus, both $\{g_{n_k}\}$ and $\{h_{n_k}\}$ are Cauchy sequences. Let
$$\lim_{k\to \infty}g_{n_k}=g,\quad \lim_{k\to \infty}h_{n_k}=h,$$
and $f=g+h$. Then, $g\in AA(X)$, $h\in M_0(X,\rho)$, $f\in PAA(X,\rho)$, and $f_{n_k}\to f$. In view of $\{f_n\}$ being a Cauchy sequence, we conclude $\lim\limits_{n\to\infty}f_n=f$. This shows that $PAA(X,\rho)$ is a Banach space under the supremum norm.
\end{proof}

By using Corollary \ref{ap}, we can get a similar result for $PAP(X,\rho)$:
\begin{corollary}For every $\rho\in \mathbb{U}_{\infty}$, $PAP(X,\rho)$ is a Banach space under the supremum norm.
\end{corollary}

\section{Acknowledgements}
The authors are grateful to the anonymous reviewers for their professional comments and suggestions, which improve greatly the quality of this paper.

\end{document}